Trace the conicoid represented by x^2 +2z^2 = y. Also describe its sections by the planes x = c,∀c ∈ R.
1
Expert's answer
2020-02-18T05:41:21-0500

It follows from the Conicoid equation "y=x^2+2\cdot z^2" that a solution exists only for "y>0". For "y<0" , there is no solution with "x,z\in R" . The form of conicoid is displayed on fig.1. For "y>0" , conicoid sections with planes "y=c\in R>0" have the form of ellipses with the formula "\frac{x^2}{c}+\frac{ z^2}{c/2}=1" (fig.2). The center of the ellipse is located on the axis Y. The small axis is directed along Z and has a length of "b=\sqrt{c/2}". The large axis of the ellipses is directed along the X axis and has a length of "a=\sqrt{c}" . The formula for this section in the new notations takes on a canonical form of ellipses "\frac{x^2}{a^2}+\frac{z^2}{b^2}=1". The conicoid body takes on values within"-a<x<a; -b<z<b".

The section by "x=c\in R" has the equation "y=c^2+2\cdot z^2" which is a parabola with an axis of symmetry parallel with axis Y. The canonical form of the parabola is "(y-c^2)=\frac{z^2}{2\cdot p}" , where "p=1\/4", and "y=c^2" is its vertex. The vertex moves out of plane XZ "(y=0)" as c increases. All the sections are simalar to each other with the focus distance of parabola "F=\frac{p}{2}=1/8". Due to symmetry about YZ plane for "c<0" we get exactly the same cross sections.

fig.1


fig.2